USE T TEST (one sample mean)


One of the undersecretary of the Department of Labor and Employment (DOLE) claims that the average salary of civil engineer is P18,000. A sample of 19 civil engineers salary has a mean of 


Given:


STEP 1: STATE THE HYPOTHESIS AND IDENTIFY THE CLAIM.


STEP 2: THE LEVEL OF SIGNIFICANCE 


STEP 3: THE T CRITICAL VALUE


STEP 4: COMPUTE THE ONE SAMPLE (T)TEST VALUE 


STEP 5: DECISION RULE


STEP 6: CONCLUSION


1
Expert's answer
2021-06-03T13:53:43-0400

Hypothesized Population Mean "\mu=18000"

Sample Standard Deviation "s=1230"

Sample Size "n=19"

Sample Mean "\bar{x}=17350"

Significance Level "\alpha=0.01"


Null and Alternative Hypotheses

The following null and alternative hypotheses need to be tested:

"H_0: \mu=18000"

"H_1: \mu\not=18000"

This corresponds to two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.01,"

"df=n-1=18" ​degrees of fredom, and the critical value for two-tailed test is"t_c=2.87844."

The rejection region for this two-tailed test is "R=\{t:|t|>2.87844\}."


The "t" - statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{17350-18000}{1230/\sqrt{19}}\approx-2.30348"

Since it is observed that "|t|=2.30348<2.87844=t_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than "18000," at the "\alpha=0.01" significance level.


Using the P-value approach: The p-value for two-tailed, the significance level "\alpha=0.01, df=18, t=-2.30348," is "p=0.033425," and since "p=0.033425>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than "18000," at the "\alpha=0.01" significance level.